AY 4 – Class Notes – Week 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

ASTRO 4: THE STARS

NOTE! This class is NOT about:

1. Planets

2. The Big Bang

3. The Expansion of the Universe

4. Galaxies

5. Dinosaur-Murdering Killer Asteroids NOTE 2! This class fulfills the “Q” requirements. “Q” = “Quantitative”. Q: If someone moved the Sun to TWICE it’s current distance, how would its brightness change?

• Qualitative answer: It would get fainter

• Quantitative answer: Based on the inverse square law with Brightness ∝ \frac{1}{Distance^{2}} ; the Sun would be only \frac{1}{2^{2}} = \frac{1}{4} as bright.

General Astronomy Questions to Get Started:

Q. Astronomy is most closely related to:

(a) Cosmetology

(b) The Human Genome Project

(c) Astrology

(d) Physics ⇐ Q. While “at” the telescope, most astronomers wear:

(a) Down jackets

(b) Trendy Patagonia Synchilla jackets

(c) Politically incorrect animal furs

(d) Bermuda shorts and loud shirts ⇐ (most observing is done from control rooms and soon it will be done completely remotely from University offices or even astronomers personal computers at home).

AY 4 – Class Notes – Week 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Some Basics Notions about the Motion of the Earth

• Q. Why is there Day and Night? The Earth rotates on its axis once every 24 hours.

• Q. Why are there Seasons? The Earth’s spin axis is tilted with respect to its orbital plane around the Sun.

The tilting of the Earth’s spin axis with respect to the plane defined by its orbit around the Sun also leads to:

• The Sun takes a different path through the sky through the year.

• Winter daylight hours are fewer and the intensity of sunlight is less.

• Those last points are the reasons why it is colder in the winter. In the Northern hemisphere we are actually a little closer to the sun in the Winter.

• Seasons are reversed in the Northern and Southern Hemispheres.

• You see different stars at different times of the year.

Moon Stuff

What time does the Full Moon rise?

(a) Midnight

(b) Noon

(c) Sunset ⇐ (See Fig 1-19 in text)

(d) Any old time and this is a trick question It takes the moon 27 days to make a trip around the Earth. Then why is it more than 29 days between full moons? (Because the Earth-Moon system is orbiting the Sun and the Moon has to travel an extra 12/360 degrees to get lined up behind the Earth a month later – Fig 1-22 in text)

Some “Q” examples

• How fast is a gaucho taking a nap at the equator traveling due to the rotation of the Earth? Earth Circumference = 24, 000 miles [π× diameter] Time for gaucho to make a “trip around the world” = 24 hours. Speed = \frac{Distance traveled}{Time taken} Speed = \frac{24,000}{24} = 1000 miles/hours

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

• How fast are we all moving through space due to the Earth’s orbit around the Sun? Distance traveled: π × 186, 000, 000 miles (diameter of the Earth’s orbit) = 584, 600, 600 miles = 5.846 × 10^{8} miles Time taken: 1 year = 365days × \frac{24hours}{1day} = 8, 760 hours Speed= \frac{5.846×10^{8}}{8.76×10^{3}} = 66, 671 miles/hour (!).

• “Dimensional Analysis”

Note that the “units” are important in these equations. They have to work out and they can be used to divine the correct equation to use if your memory or physical reasoning abilities abandon you for some reason. For example: Q. How fast are we moving through... Suppose you knew this was somehow related to how far we travel and how long it takes but you can’t remember the exact equation. You CAN remember that “how fast” must come out in units of velocity like Miles/Hour. So, lets try:

Speed = Distancetraveled × Timetaken

The units for the multiplication on the right-hand side are

Distance-time -> miles-hour

This can’t be right. So, work backwards from the units, the way to end up with the right velocity units is clearly to take the distance traveled divided by the time taken. Then the units are:

Distance/Time -> mile/hour

Telescopes

Telescopes and the equipment that go long with them are very analogous to human eyes and brains. The advantage of telescopes is they have considerably more light gathering area. Telescopes have to do (only) a few things:

• Focus the light from objects onto a detector

• Point to various places on the sky

• “Track” objects as they drift across the sky

At the end of the last century, the biggest telescopes had mirrors with diameters of ∼ 40 inches (1 meter) and observations were “visual”. There was a steady increase in mirror size and mechanical sophistication up through the 1950’s when the Palomar 200” (5 meter) telescope went into operation. The Soviets built a 6m mirror telescope that never worked very well. The Palomar reigned as the biggest useful telescope will 1994 when the first of two Keck telescopes was completed which had a 10m (400”, 33.3 foot) diamter mirror.

• How much more light does the Keck 10m telescope gather than your eye with it light collecting diameter of 0.5 inch? Ratio of light gathering AREA:

(\frac{400}{0.5})^{2} = 600, 000

Note: used same units for Keck and eye diameters It is hard to get a feel for these big

numbers sometimes. A 2 candlepower light on the moon would give 9 photons per minute with Keck and 1 photon per 46 days with your naked eye. This is why we build big telescopes. Go outside on a dark night an look up. There are about 3000 stars visible to the

“naked eye” and 3 galaxies. With the Keck telescope, there are billions of stars visible and billions of galaxies.

• Where are these telescopes?

– Mt Hamilton (3m and 1m)

– Kitt Peak National Observatory (Arizona) and the

– Cerro Tololo Interamerican Observatory (Chile) are the main US national facilities (4m, 2m, 1m)

– On top of Mauna Kea on Hawaii there is a whole community of telescopes including the 2 Keck 10m telescopes, an 8m Japanese telescope, radio telescopes, the Canada-France-Hawaii telescopes and soon, an 8m national facility telescope.

– ESO telescopes, Canary Islands, Australia

• Why always put telescopes on top of mountains?

• The Hubble Space Telescope – higher than your average mountaintop

(1) HST gives very high spatial resolution

(2) HST has “darker” skies

(3) HST opens up new spectral windows (more later)

• What is the US budget for astronomy? ∼ $200 million/year = 65 cents per year per

person. One Keck telescope runs about $80 million to build and $10 million to run.